# Enabling Efficient and Flexible Interpretability of Data-driven Anomaly Detection in Industrial Processes with AcME-AD

### v1

@@ V Conclusions
This work successfully demonstrates the application of our proposed method, AcME-AD, for anomaly detection in an industrial setting. Firstly, AcME-AD facilitates user decision-making through clear and informative visualizations. The integrated ”what-if” tool for local interpretation empowers users to pinpoint the root cause of anomalies and take appropriate corrective actions. Secondly, its efficient computation enables seamless integration into decision support systems, promoting real-time anomaly detection and response. We highlight the particular suitability of AcME-AD for unsupervised anomaly detection tasks, which are of paramount importance in industrial applications. By providing model-agnostic explanations, AcME-AD empowers domain experts to select models that best align with their knowledge, fostering crucial human-machine collaboration in the context of Industry 5.0. As future development of the present work, we aim at refining the visualizations based on feedback from real-world deployments. Additionally, we plan to explore the potential of utilizing local explanations for clustering anomalies into distinct groups, thereby enabling supervised classification tasks.
@@ Acknowledgments
This work was partially carried out within the MICS (Made in Italy – Circular and Sustainable) Extended Partnership and received funding from Next-GenerationEU (Italian PNRR – M4C2, Invest 1.3 – D.D. 1551.11-10-2022, PE00000004). Moreover this study was also partially carried out within the PNRR research activities of the consortium iNEST (Interconnected North-Est Innovation Ecosystem) funded by the European Union Next-GenerationEU (Piano Nazionale di Ripresa e Resilienza (PNRR) – Missione 4 Componente 2, Investimento 1.5 – D.D. 1058 23/06/2022, ECS00000043). This work was also cofunded by the European Union in the context of the Horizon Europe project ‘AIMS5.0 - Artificial Intelligence in Manufacturing leading to Sustainability and Industry5.0’ Grant agreement ID: 101112089.

### v2

@@ V Conclusions
This work successfully demonstrates the application of our proposed method, AcME-AD, for anomaly detection in an industrial setting.
Firstly, AcME-AD facilitates user decision-making through clear and informative visualizations. The integrated ”what-if” tool for local interpretation empowers users to pinpoint the root cause of anomalies and take appropriate corrective actions. Secondly, its efficient computation enables seamless integration into decision support systems, promoting real-time anomaly detection and response.
We highlight the particular suitability of AcME-AD for unsupervised anomaly detection tasks, which are of paramount importance in industrial applications. By providing model-agnostic explanations, AcME-AD empowers domain experts to select models that best align with their knowledge, fostering crucial human-machine collaboration in the context of Industry 5.0. As future development of the present work, we aim at refining the visualizations based on feedback from real-world deployments. Additionally, we plan to explore the potential of utilizing local explanations for clustering anomalies into distinct groups, thereby enabling supervised classification tasks.
@@ Acknowledgments
This work was partially carried out within the MICS (Made in Italy – Circular and Sustainable) Extended Partnership and received funding from Next-GenerationEU (Italian PNRR – M4C2, Invest 1.3 – D.D. 1551.11-10-2022, PE00000004). Moreover this study was also partially carried out within the PNRR research activities of the consortium iNEST (Interconnected North-Est Innovation Ecosystem) funded by the European Union Next-GenerationEU (Piano Nazionale di Ripresa e Resilienza (PNRR) – Missione 4 Componente 2, Investimento 1.5 – D.D. 1058 23/06/2022, ECS00000043). This work was also cofunded by the European Union in the context of the Horizon Europe project ‘AIMS5.0 - Artificial Intelligence in Manufacturing leading to Sustainability and Industry5.0’ Grant agreement ID: 101112089. The Regione Veneto is also gratefully acknowledged for co-financing part of this activity through the Development and Cohesion Plan (PSC) under the special section THEME AREA PSC 1 - RESEARCH AND INNOVATION, in continuity with ACTION (POR FESR) 1.1.4 ’Support for collaborative R&D activities for the development of new sustainable technologies, products, and services,’ as per DGR no. 1800 dated December 15, 2021 - Project MIMIC.